# What should a neuron aim for? Designing local objective functions based on information theory

### v3

@@ A.1 Pseudocode for training algorithm (Setup 1)
In this Section, we provide pseudocode detailing the procedure to train infomorphic networks with dense lateral connections and label context (i.e., Setup 1 in Figure 3). The code is started from the main function TrainModel, which invokes TrainNeuronTrivariate and TrainNeuronBivariate to update hidden layer and output layer weights, respectively. These functions in turn make use of the auxiliary functions ComputeIsxRedundancies, which computes generalized redundancies according to the analytical definition by [14], and ComputePIDAtoms, which performs a Moebius inversion to compute the PID atoms from the generalized redundancies [12]. Note that since the linear operation performed by ComputePIDAtoms is fixed, it can be efficiently implemented as a multiplication with a precomputed matrix.
Note that, as originally proposed by [18] and utilized by [17], gradients for the incoming weights are only computed with respect to the conditional probability $p(y|f,c,l)$ (or $p(y|f,c)$) of the output $Y$, since the empirical probability mass functions $p(f,c,l)$ in Algorithm 2 and $p(f,c)$ in Algorithm 3 are not differentiable. While we currently have no strong a priori argument why the empirical probability mass functions can be assumed constant, the learning of PID over training and the good task performance when using these gradients in experiments provides an ex post justification for this procedure.
@@ 3 COMPUTE empirical probability masses p(f, c);
4 EVALUATE conditional probabilities p(y|f, c, l) from the neurons;
@@ A.2 Definition of shared-exclusion redundancy
In this section, we briefly motivate and explain the shared-exclusion redundancy measure $I_{\cap}^{\mathrm{sx}}$ introduced by Makkeh et al. [14]. The mutual information $I(T:S_{1},S_{2})$ between a target variable $T$ and two source variables $S_{1}$ and $S_{2}$ can be interpreted in a Bayesian sense as an average measure for how the prior belief of the target event $T=t$, needs to be updated in light of the event of observing both source events $S_{1}=s_{1}$ and $S_{2}=s_{2}$ simultaneously:
@@
Makkeh et al. [14] build on this logic and define redundancy as an average measure for how the beliefs about the target event $T=t$ need to be updated if instead it is only known that $S_{1}=s_{1}$ or $S_{2}=s_{2}$ have occurred:
@@
For more than two source variables $s_{i}$ (where i is an index enumerating the set of source variables), the term to condition on becomes a disjunction between conjunctions of the form $\bigvee_{\bm{a}\in\alpha}\bigwedge_{i\in\bm{a}}S_{i}=s_{i}$ for the redundancy associated with the antichain $\alpha$. The atoms $\Pi$ can then be computed from these redundancies via a Moebius inversion, as laid out in detail in Gutknecht et al. [13].
The definition is symmetric with respect to permutation of the sources, fulfills a target chain rule and is differentiable with respect to the underlying probability distribution [14], which makes it a suitable definition for optimizing objective functions.
@@ A.3 Compute resources
For the figures in this paper, we performed a total of seven hyperparameter optimizations: Three TPE optimizations and three CMA-ES optimizations of MNIST objective function parameter and a single CMA-ES optimization for the CIFAR task. Each of these optimizations took $\leq 12$ hours to compute on a HPC cluster node with $32$ cpu cores and two NVIDIA A100 GPUs with 404040\,40GB of VRAM each.
@@ A.4 Implementation details and model parameters
We implemented infomorphic networks as a flexible and efficient python package using pytorch [53] for automatic differentiation of the local goal functions. Furthermore, the optuna [54] package was used to compute the TPE and CMA-ES hyperparameter optimizations as well as the computation of the parameter importance via the mean decrease in impurity.
In this paper, we use the MNIST [55] and CIFAR10 [22] datasets. The MNIST dataset consists of 70,000 grayscale images of handwritten digits, each sized 28x28 pixels. The dataset is split into a training set of 60,000 samples and a test set of 10,000 samples. The CIFAR10 dataset consists of 60,000 images with three color channels and a resolution of 32x32 pixels. Of the 60,000 images, 50,000 are in the training set and 10,000 are in the test set. For each of our training runs with either dataset, $20\%$ of the training set samples are withheld randomly to be used for validation.
@@
Table 1: Model and training parameters for the models shown in Fig. 3. *The effect of different batch sizes is analyzed in Fig. 10. **The efffect of the number of bins is analyzed in Fig. 11
@@ A.5 Model Dynamics During Training
where $\boldsymbol{w}_{F}$ corresponds to feedforward weights of a single neuron of the hidden layer. The median value of $D_{c}^{(t)}$ for the hidden neurons of a trivariate model are shown in Fig. 5. While the self-cosine distance consistently increases with layer size in the dense setups, it does not increase in the sparse setup after a layer size of 100 neurons which is also the number of lateral connections in all of the larger sparse layers.
@@ A.6 Optimized Objective Functions for the Trivariate Neurons
To support the trivariate objective function that was obtained using hyperparameter optimization, we repeated the optimizations three times for both, the TPE sampler and the CMA-ES sampler three times each. The resulting objective functions are illustrated in Fig. 6. While the objective functions that were optimized with the TPE sampler included all $\gamma$ parameters, the optimization with the CMA-ES sampler was only performed for the parameters that describe PID-atoms while the parameter for the residual entropy was set to 0. This was done because we observed a strong correlation between the parameter for the unique information of the feedforward input and the residual entropy which both had to be small. In the figure, the respective values are set 0. In the figure, we also illustrate the mean decrease impurity score which is a measure for the importance of a parameter.
Figure 6: The three objective functions shown in the first column were obtained by utilizing the TPE sampler while the objective functions in the third column were obtained with the CMA-ES sampler. The performance illustrated below each of the objective functions corresponds to the maximum validation accuracy during optimization. The column next to the objective functions show the corresponding mean decrease impurity score of the corresponding goal parameter as a measure for the importance of the parameters. While the objective functions obtained with CMA-ES have a higher maximum validation accuracy than the goals obtained with the TPE, we observed that the first objective function obtained with the TPE sampler (top-left) outperformed the others in terms of median validation accuracy for larger layer sizes, which is why we used this objective function for the main results.
@@ A.7 Trivariate Consistency Equations
A.8 The Atoms of the Partial Information Decomposition for Three Source Variables
Figure 7: To improve the clarity and intuition for the atoms of the partial information decomposition with three source variables, this figure shows a larger version of Venn-diagram of the decomposition already shown in Fig. 2.C. The meaning of the different information atoms are also listed in Table 2 as a guide for the reader.
@@ A.8 The Atoms of the Partial Information Decomposition for Three Source Variables
$\Pi_{\{FC\}\{FL\}}$ information provided by both, the synergy between the feedforward and the context input and the synergy between the feedforward and the context input
@@
Table 3: The four most important non-zero goal parameters found through optimization on MNIST and ablation lead to interpretable requirements on each neuron’s local information processing.
@@
$\gamma_{\{F\}\{C\}}$ redundancy between $F$ and $C$ 0.98 maximize information that is provided both by feedforward AND context but not by other neurons – and thus relevant for the task
$\gamma_{\{F\}\{L\}}$ redundancy between $F$ and $L$0.99 minimize information that is redundant with other neurons, but NOT provided by the context – and thus not relevant for the task
$\gamma_{\{F\}\{C\}\{L\}}$ redundancy between $F$, $C$ and $L$ 0.33 moderately maximize information that is redundant with other neurons AND the context – thus relevant, but already encoded
@@ A.9 Recurrence of the Hidden Layer
To perform an information decomposition which includes the information of the lateral activity, we include lateral connections within the hidden layer as illustrated in Fig. 3.A. As described in Section A.1, we only perform two iterations within the hidden layer i.e. one iteration to compute the activity of the neurons within the layer and a second iteration to include this activity in the final output of the neurons. Therefore, the lateral activity contributes to the neurons output. Due to the stochasticity of the neurons, the lateral input will never fully converge. The continuous output probabilities converge already after the second single iteration with only minor changes in the output probabilities of the order of $10^{-4}$. This is due to the small influence of the lateral input on the output which is limited to $0.1\,\sigma(2FL)$ i.e. a maximum change of ±$\pm 0.1$ in the activation.
@@ A.10 Deep infomorphic networks
A major factor in the success in many neural network architectures is their layered structure, allowing the network to encode the relevant information in successive steps to make them more accessible to the final output layer. For this reason, scaling infomorphic networks to multiple hidden layers is the canonical next step for making them applicable to more complex tasks.
@@
As shown in figure 2.D we now created the goal function for the trivariate neurons by maximizing the atom that corresponds to the overlap of the two bivariate atoms. This atom corresponds to the redundant information between the feedforward and the contextual input, that is uniquely provided by the two sources and not by the lateral input, resulting in the heuristic goal function of the form $G_{\text{triv}}=\Pi_{\{F\}\{C\}}$. It is important to note, that even if this goal looks identical to $G_{\text{con}}$, it is different as the meaning of the atoms depends on the number of sources despite the identical notation.
@@ A.11 Batch Size and Number of Bins
Figure 10: To evaluate whether a batch size of 1024 is sufficient for estimating the probability distributions required for information decomposition, we conducted additional experiments with the trivariate model. These experiments utilized the optimized goal function, the parameters listed in Table 1, and a hidden layer size of $N_{\text{hid}}=100$, while varying the batch size. Each configuration was run in 10 different initializations and trained for 1000 epochs to ensure convergence, as demonstrated in the right figure. The left figure reveals that performance remains consistent across different batch sizes. However, models with smaller batch sizes converge faster, reducing the computational cost of optimization. Consequently, a batch size of 1024 appears adequate for the requirements of this study.

### v4

@@ A.1 Pseudocode for Training Algorithm (Setup 1)
In this section, we provide pseudocode detailing the procedure to train infomorphic networks with dense lateral connections and label context (i.e., Setup 1 in Figure 3). The code is started from the main function TrainModel, which invokes TrainNeuronTrivariate and TrainNeuronBivariate to update hidden layer and output layer weights, respectively. These functions in turn make use of the auxiliary functions ComputeIsxRedundancies, which computes generalized redundancies according to the analytical definition by Makkeh et al. (2021), and ComputePIDAtoms, which performs a Moebius inversion to compute the PID atoms from the generalized redundancies (Williams and Beer, 2010). Note that since the linear operation performed by ComputePIDAtoms is fixed, it can be efficiently implemented as a multiplication with a precomputed matrix.
Note that, as originally proposed by Kay (1994) and utilized by Makkeh et al. (2025), gradients for the incoming weights are only computed with respect to the conditional probability $p(y|f,c,l)$ (or $p(y|f,c)$) of the output $Y$, since the empirical probability mass functions $p(f,c,l)$ in Algorithm 2 and $p(f,c)$ in Algorithm 3 are not differentiable. While we currently have no strong a priori argument why the empirical probability mass functions can be assumed constant, the learning of PID over training and the good task performance when using these gradients in experiments provides an ex post justification for this procedure.
@@ 3 COMPUTE empirical probability masses p(f, c);
4 EVALUATE conditional probabilities p(y|f, c) from the neurons;
@@ A.2 Definition of Shared-Exclusion Redundancy
In this section, we briefly motivate and explain the shared-exclusion redundancy measure $I_{\cap}^{\mathrm{sx}}$ introduced by Makkeh et al. (2021). The mutual information $I(T:S_{1},S_{2})$ between a target variable $T$ and two source variables $S_{1}$ and $S_{2}$ can be interpreted in a Bayesian sense as an average measure for how the prior belief of the target event $T=t$, needs to be updated in light of the event of observing both source events $S_{1}=s_{1}$ and $S_{2}=s_{2}$ simultaneously:
@@
Makkeh et al. (2021) build on this logic and define redundancy as an average measure for how the beliefs about the target event $T=t$ need to be updated if instead it is only known that $S_{1}=s_{1}$ or $S_{2}=s_{2}$ have occurred:
@@
For more than two source variables $s_{i}$ (where i is an index enumerating the set of source variables), the term to condition on becomes a disjunction between conjunctions of the form $\bigvee_{\bm{a}\in\alpha}\bigwedge_{i\in\bm{a}}S_{i}=s_{i}$ for the redundancy associated with the antichain $\alpha$. The atoms $\Pi$ can then be computed from these redundancies via a Moebius inversion, as laid out in detail in Gutknecht et al. (2021).
The definition is symmetric with respect to permutation of the sources, fulfills a target chain rule and is differentiable with respect to the underlying probability distribution (Makkeh et al., 2021), which makes it a suitable definition for optimizing objective functions.
@@ A.3 Compute Resources
For the figures in this paper, we performed a total of seven hyperparameter optimizations: Three TPE optimizations and three CMA-ES optimizations of MNIST objective function parameter and a single CMA-ES optimization for the CIFAR-10 task. Each of these optimizations took $\leq 12$ hours to compute on a HPC cluster node with $32$ cpu cores and two NVIDIA A100 GPUs with 404040\,40GB of VRAM each.
@@ A.4 Implementation Details and Model Parameters
We implemented infomorphic networks as a flexible and efficient python package using pytorch (Paszke et al., 2019) for automatic differentiation of the local goal functions. Furthermore, the optuna (Akiba et al., 2019) package was used to compute the TPE and CMA-ES hyperparameter optimizations as well as the computation of the parameter importance via the mean decrease in impurity.
In this paper, we use the MNIST (LeCun et al., 2015) and CIFAR-10 (Krizhevsky, 2012) datasets. The MNIST dataset consists of 70,000 grayscale images of handwritten digits, each sized 28x28 pixels. The dataset is split into a training set of 60,000 samples and a test set of 10,000 samples. The CIFAR-10 dataset consists of 60,000 images with three color channels and a resolution of 32x32 pixels. Of the 60,000 images, 50,000 are in the training set and 10,000 are in the test set. For each of our training runs with either dataset, $20\%$ of the training set samples are withheld randomly to be used for validation.
@@
Table 1: Model and training parameters for the models shown in Fig. 3. *The effect of different batch sizes is analyzed in Fig. 10. **The effect of the number of bins is analyzed in Fig. 11
@@ A.5 Model Dynamics During Training
where $\boldsymbol{w}_{F}$ corresponds to feedforward weights of a single neuron of the hidden layer. The median value of $D_{c}^{(t)}$ for the hidden neurons of a trivariate model are shown in Fig. 5. While the self-cosine distance consistently increases with layer size in the dense setups, it does not increase in the sparse setup after a layer size of 100 neurons, which is also the number of lateral connections in all of the larger sparse layers.
@@ A.6 Optimized Objective Functions for the Trivariate Neurons
To support the trivariate objective function that was obtained using hyperparameter optimization, we repeated the optimizations three times for both, the TPE sampler and the CMA-ES sampler, with parameter values being limited to the interval $\gamma_{i}\in[-1,1]$. The resulting objective functions are illustrated in Fig. 6. While the objective functions that were optimized with the TPE sampler included all $\gamma$ parameters, the optimization with the CMA-ES sampler was only performed for the parameters that describe PID-atoms while the parameter for the residual entropy was set to 0. This was done because we observed a strong correlation between the parameter for the unique information of the feedforward input and the residual entropy which both had to be small. In the figure, the respective values are set 0. In the figure, we also illustrate the mean decrease impurity score which is a measure for the importance of a parameter.
Figure 6: The three objective functions shown in the first column were obtained by utilizing the TPE sampler while the objective functions in the third column were obtained with the CMA-ES sampler. The performance illustrated below each of the objective functions corresponds to the maximum validation accuracy during optimization. The column next to the objective functions shows the corresponding mean decrease impurity score of the corresponding goal parameter as a measure for the importance of the parameters. While the objective functions obtained with CMA-ES have a higher maximum validation accuracy than the goals obtained with the TPE, we observed that the first objective function obtained with the TPE sampler (top-left) outperformed the others in terms of median validation accuracy for larger layer sizes, which is why we used this objective function for the main results.
@@ A.7 Trivariate Consistency Equations
A.8 The Atoms of the Partial Information Decomposition for Three Source Variables and their Interpretation
To improve the clarity and intuition for the atoms of the partial information decomposition with three source variables, Fig. 7 shows a larger version of decomposition diagram already shown in Fig. 2.C. As a guide for the reader, the meaning of the different information atoms are listed in Table 2. Additionally, Table 3 explicitly lists the goal parameter values found via optimization together with their interpretation.
Figure 7: Larger version of decomposition diagram already shown in Fig. 2.
@@
$\Pi_{\{FC\}\{FL\}}$ information provided by both, the synergy between the feedforward and the context input and the synergy between the feedforward and the lateral input
@@
Table 3: The four most important non-zero goal parameters found through optimization on MNIST and ablation lead to interpretable requirements on each neuron’s local information processing. Note that the optimization was performed with parameter values being limited to the interval $\gamma_{i}\in[-1,1]$.
@@
$\gamma_{\{F\}\{C\}}$ redundancy between $F$ and $C$ 0.98 maximize information that is provided both by feedforward AND context but not by other neurons—and thus relevant for the task
$\gamma_{\{F\}\{L\}}$ redundancy between $F$ and $L$0.99 minimize information that is redundant with other neurons, but NOT provided by the context—and thus not relevant for the task
$\gamma_{\{F\}\{C\}\{L\}}$ redundancy between $F$, $C$ and $L$ 0.33 moderately maximize information that is redundant with other neurons AND the context—thus relevant, but already encoded
@@ A.9 Recurrence of the Hidden Layer
To perform an information decomposition which includes the information of the lateral activity, we include lateral connections within the hidden layer as illustrated in Fig. 3.A. As described in Section A.1, we only perform two iterations within the hidden layer i.e.,one iteration to compute the activity of the neurons within the layer and a second iteration to include this activity in the final output of the neurons. Therefore, the lateral activity contributes to the neurons output. Due to the stochasticity of the neurons, the lateral input will never fully converge. The continuous output probabilities converge already after the second single iteration with only minor changes in the output probabilities of the order of $10^{-4}$. This is due to the small influence of the lateral input on the output which is limited to $0.1\,\sigma(2FL)$ i.e., a maximum change of ±$\pm 0.1$ in the activation.
@@ A.10 Deep Infomorphic Networks
A major factor in the success in many neural network architectures is their layered structure, allowing the network to process the relevant information in successive steps to make them more accessible to the final output layer. For this reason, scaling infomorphic networks to multiple hidden layers is the canonical next step for making them applicable to more complex tasks.
@@
As shown in figure 2.D we now created the goal function for the trivariate neurons by maximizing the atom that corresponds to the overlap of the two bivariate atoms. This atom corresponds to the redundant information between the feedforward and the contextual input, that is uniquely provided by the two sources and not by the lateral input, resulting in the heuristic goal function of the form $G_{\text{triv}}=\Pi_{\{F\}\{C\}}$. It is important to note, that even if this goal looks identical to the bivariate goal that maximizes the redundant information between F and C, it is different as the meaning of the atoms depends on the number of sources despite the identical notation.
@@ A.11 Batch Size and Number of Bins
Figure 10: To evaluate whether a batch size of 1024 is sufficient for estimating the probability distributions required for information decomposition, we conducted additional experiments with the trivariate model. These experiments utilized the optimized goal function, and a hidden layer size of $N_{\text{hid}}=100$, while varying the batch size. Each configuration was run in 10 different initializations and trained for 1000 epochs to ensure convergence, as demonstrated in the right figure. The left figure reveals that performance remains consistent across different batch sizes. However, models with smaller batch sizes converge faster, reducing the computational cost of optimization. Consequently, a batch size of 1024 appears adequate for the requirements of this study.